# CoDial: Interpretable Task-Oriented Dialogue Systems Through Dialogue Flow Alignment

### v1

@@ 3.3 CoDial Human Feedback Integration
CoDial’s Human Feedback (CHF) mechanism incorporates feedback to refine or improve the generated dialogue logic. The code enhancement through human feedback comprises two broad approaches: i) manual modifications and ii) LLM-aided modifications.
CHF assist human feedback incorporation in the form of refinement instructions (RIs), shown at the top in Figure 1. RIs allow the domain expert or developer of CoDial to refine the generated logic through text-to-code instructions. As the first step, we provide three instructions for refining certain aspects of the output code: correct logic (i.e., if statement) for each node, DST initialization, and request node checks. These RIs, presented in Table 6, are always applied to fix the problems in the output, if any. Moreover, we attempt to refine the generated DST prompts ($P^{(s)}$) through automatic prompt optimization. Please refer to Section A.3 for details on automatic DST prompt optimization. In addition, CHF allows for manual modifications on the dialogue flow (Section A.2) and manual DST prompt optimization (Section A.3).
@@ 4 Experimental Settings
We use GPT-4o [^5] https://openai.com/index/hello-gpt-4o/, Claude 3.5 Sonnet [^6] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 2.0 Flash [^7] https://developers.googleblog.com/en/gemini-2-family-expands/, and DeepSeek V3 (DSV3) (DeepSeek-AI et al., 2024) as $\text{LLM}_{\text{GCG}}$, and GPT-4o-mini and DSV3 as $\text{LLM}_{\text{A}}$ [^8] All models were accessed from January to May 2025. . Larger models are used for code generation—given the complexity of the task, we found that smaller models often fail to fully adhere to instructions. For further details, please refer to Section A.4.
@@ 4.1 Datasets
The STAR dataset (Mosig et al., 2020), collected in a Wizard-of-Oz setup (human-human conversations), provides explicit task schemas (i.e., dialogue flows) to ensure consistent and deterministic system actions. We also use silver state annotations created in STARv2 (Zhao et al., 2023) for ablation studies. Refer to Section A.2 for more implementation details on STAR.
@@
MultiWOZ (Budzianowski et al., 2018) is a large-scale, multi-domain TOD dataset consisting of human-human conversations, with most domains involving booking subtasks such as hotel reservations and taxi services. Given the impracticality of crafting dialogue flows for every possible domain combination (Zhang et al., 2023), we report results in a naive oracle domain setting. Please refer to Section A.3 for more details.
@@ 4.2 Metrics
For the STAR dataset, we compute BLEU-4 score (Papineni et al., 2002). We also follow Mosig et al. (2020) to compute F1 and accuracy. For the MultiWOZ dataset, we compute BLEU, Inform and Success rates, and Joint Goal Accuracy (JGA) using the official evaluation script (Nekvinda and Dušek, 2021). We report the mean result of three runs. Please refer to Section A.4 for more implementation details.
@@ 4.3 Baselines
For a complete list of compared methods, please refer to Section A.5. Our most comparable baselines are as follows:
@@
  • SGP-TOD (Zhang et al., 2023) is a zero-shot prompting approach that employs graph-based dialogue flows and out-of-domain formatting examples. Refer to Section A.5 for details on fair comparison.
@@ 5 Experimental Results
5.1 Results and Analysis on STAR
Model Approach $\text{LLM}_{\text{GCG}}$ $\text{LLM}_{\text{A}}$ RI F1 Acc. BLEU
Training-based Approaches Domain Transfer
BERT + Schema Fine-tuning - - - 29.7 32.4 -
SAM Fine-tuning - - - 51.2 49.8 -
AnyTOD norec base Fine-tuning - - - 55.8 56.1 32.4
AnyTOD prog+sgd xxl Pretraining - - - 70.7 70.8 44.2
Inference-based Approaches Strict Zero-shot
SGP-TOD-gpt3.5-e2e Zero-shot - - - 53.5 53.2 -
CoDial (Ours)
CoDial original DFs Zero-shot 4o 4o-mini 51.9 51.1 38.9
CoDial -- ri Zero-shot 4o 4o-mini 56.1 57.3 38.4
CoDial -- ri Zero-shot Sonnet 4o-mini 57.0 58.4 39.2
CoDial Zero-shot DSV3 4o-mini 46.1 48.0 28.0
CoDial* Zero-shot Gem. 2 Fl. 4o-mini 50.5 52.1 32.9
CoDial Zero-shot Sonnet 4o-mini 57.7 58.5 39.3
CoDial Zero-shot 4o DSV3 55.6 56.8 44.2
CoDial Zero-shot 4o 4o-mini 58.5 60.1 45.2
@@ 5.1 Results and Analysis on STAR
Table 1: Comparison of models on the STAR dataset. Our CoDial model achieves the SOTA in a “Strict Zero-Shot” setting, where we do not require any training samples or sample conversations. SAM results are cited from Zhao et al. (2023). The generated code for the model with an asterisk (*) has been manually fixed and is not directly comparable. DF stands for “dialogue flow.”
SOTA Performance Without Training.
Table 1 summarizes our results on the STAR dataset. CoDial achieves strong performance, surpassing all training-based approaches except AnyTOD prog+sgd xxl, and sets the new SOTA among inference-based methods. Our framework improves F1 by +5 and accuracy by +6.9 points over the previous SOTA. While AnyTOD prog+sgd xxl achieves higher scores, it requires manually written dialogue logic programs, making it less accessible to non-programmers, and extensive pre-training with task-specific data. In contrast, CoDial operates in a strict zero-shot setting, eliminating the need for manual programming and training. Without modifications (Section A.2), the original STAR dialogue flows result in lower performance (F1: $51.9$). After manually modifying the dialogue flow and applying LLM-aided modifications to the generated code, we significantly enhance performance. We further explore the impact of LLM-aided corrections in Section 5.3.
Impact of Model Selection
@@
We experience with different model choices for the ($\text{LLM}_{\text{GCG}}$, $\text{LLM}_{\text{A}}$) pairing. Better instruction following and more robust code generation often translate to higher overall performance. Because most LLMs are unfamiliar with guardrailing languages such as Colang, they must accurately interpret the $\text{prompt}_{\text{GCG}}$ to produce syntactically correct code. When the chosen LLM struggles with instruction following, code generation can fail, leading to incorrect or incomplete programs. Among the tested configurations, CoDial (4o, 4o-mini) achieves the highest performance in all metrics. We also report results in an oracle voting setting (Table 5) between GPT-4o-mini and DSV3 as $\text{LLM}_{\text{A}}$, where for each task, we take the best-performing $\text{LLM}_{\text{A}}$ by F1. This results in an increase of $+1.7$ F1 and $+1.5$ accuracy.
State and Action Prediction and API Calls
We find that NeMo Guardrails’ intent detection performs strongly, achieving an F1 score of 96.3 on global actions (Table 5). Additionally, we observe that STAR’s API calling precision—measured as the ratio of correct API calls to the total number of API calls—stands at 74.9. Table 5 also summarizes the performance of the actions that are generated by $\text{LLM}_{\text{A}}$ (i.e., when NAP component does not generate an output). LLM-generated actions account for $25\%$ of all predicted actions, with $70\%$ of them belonging to three fallback actions: goodbye, out_of_scope, and anything_else. Excluding fallbacks, LLM-generated actions only account for $9.2\%$ of predictions, indicating that our NAP logic is generally effective at generating outputs based on the predicted state. Since fallback actions are a simple 3-way classification, we would expect high performance. However, $\text{LLM}_{\text{A}}$ achieves an F1 score of only $51.4$. We attribute this to the lack of an explicit schema for fallback actions in the STAR dataset, leading to inconsistencies in wizard annotations. Additionally, we observe a significant performance drop from fallback to non-fallback actions in both F1 ($51.4\rightarrow 38.7$) and accuracy. This suggests that despite having an explicit schema, LLMs struggle to capture the more complex logic needed to predict non-fallback actions. Our findings align with Dong et al. (2024), reinforcing the need for a neuro-symbolic approach.
Figure 3 shows the error rate of the predicted conversation state across different node types for each model. To approximate the error, we compare the model’s predicted state with the estimated ground-truth state (i.e., the wizard’s state), as described in Section A.2. We find that the error rate generally inversely correlates with the overall performance in Table 1; higher-performing models tend to exhibit lower state prediction error.
Model JGA Inform Success BLEU Combined
Training-based Approaches
SOLOIST 35.9 81.7 67.1 13.6 88.0
MARS 35.5 88.9 78.0 19.6 103.0
AnyTOD XXL 30.8 76.9 47.6 3.4 65.6
Inference-based Approaches
IG-TOD (fs) 27 - 44 6.8 -
SGP-TOD - 82.0 72.5 9.2 86.5
CoDial 28.4 76.6 54.6 3.5 69.1
Table 2: Comparison of models on the MultiWOZ dataset. SOLOIST and MARS results are cited from Zhao et al. (2023). (fs) indicates few-shot.
@@
Table 5: Ablations on MultiWOZ. Settings with an asterisk (*) are not directly comparable due to a simpler task setup. (fs) indicates few-shot.
5.2 Results on MultiWOZ
Our results on the MultiWOZ dataset are summarized in Table 2. Unlike STAR, where wizards were provided structured guidance for system responses, MultiWOZ lacks a predefined dialogue flow, making interactions less consistent. This variability in MultiWOZ poses additional challenges for heuristics-grounded and programmatic approaches like CoDial and our most comparable system, AnyTOD. AnyTOD trains LLMs to use a strict programmatic schema and generates output text with templates. However, we achieve comparable performance to AnyTOD without pre-training.
Most of the MultiWOZ test set consists of multi-domain conversations, where a user may, for example, book both a taxi and a restaurant in the same dialogue. Since CoDial is designed for single-domain interactions, we report its performance on single-domain dialogues in Table 5, where it performs well. However, with our naive oracle domain setting, CoDial performance drops significantly. This is likely due to compounded errors from DST to NAP, which we analyze further in Section 5.3.
5.3 Detailed Analysis
@@
To assess the impact of compounded DST error, we evaluate CoDial under an Oracle setting. Since STAR does not provide gold DST labels, we simulate an oracle setting by replacing CoDial’s DST component with the silver-annotations from STARv2 (Table 5). This results in a performance gain of $+2.2$ F1 and $+2.8$ accuracy. We do the same for MultiWOZ, where we replace our predictions with the gold belief state. As shown in Table 5, this leads to a substantial performance improvement, making CoDial competitive with other baselines. These findings suggest that exploring more advanced DST approaches could be a promising direction to improve performance, regardless of the dataset. We experiment briefly with prompt optimization, described below, but we urge further exploration in this direction.
@@
We use LLMs to perform iterative code refinement and automatic prompt optimization for the DST prompts. Refining the code with RIs consistently enhances CoDial’s performance, demonstrating the benefits of integrating user feedback into the generation process. After prompting the LLM to iteratively refine its outputs, CoDial achieves better accuracy and fluency (compared to CoDial -- ri in Table 1). We also conduct an ablation study to examine the effect of the individual RIs, summarized in Table 5. Although all RIs are beneficial, most of the performance improvements can be attributed to the third RI, which refines the conditional logic of request nodes.
After observing the results of the oracle DST setting, we also apply prompt optimization to improve DST accuracy. As shown in Table 5, automatic prompt optimization yields only marginal gains across metrics, with the exception of Inform, suggesting that improving DST remains a non-trivial challenge—potentially due to the limitations of our naive multi-domain evaluation setup. This is a surprising result, but it suggests that not all DST slots are equally valuable for the final performance. To explore the impact of human feedback, we also experiment with manual prompt optimization (Section A.3), making minor edits to the prompts for the “attraction” domain. This results in consistent improvements across all metrics, reinforcing that human-crafted prompts can still outperform automatic optimization.
@@
To better understand the effectiveness of the proposed CoDial architecture, we experiment with a setting in which the NAP component is removed and all actions are predicted in a fully generative manner by the $\text{LLM}_{\text{A}}$, similar to Zhang et al. (2023). We prompt the $\text{LLM}_{\text{A}}$ with simplified dialogue flows following their work and include predicted DST slots in the prompt. As shown in Table 5, this results in a substantial drop in performance, highlighting the importance of our structured NAP logical flow approach. We further ablate the model by removing the DST component, causing a larger drop as expected.

### v2

@@ 3.3 CoDial Human Feedback Integration
CoDial’s Human Feedback (CHF) mechanism incorporates human feedback to refine or improve the generated guardrailing code $g$. The code enhancement through feedback comprises two broad approaches: i) manual and ii) LLM-aided modifications.
CHF assists iterative improvement of $g$ in the form of refinement instructions (RIs), shown at the top in Figure˜1. RIs allow the user of CoDial to refine the generated logic through hand-crafted instructions. We provide three instructions for refining certain aspects of the output code: correct logic (i.e., if statement) for each node, DST initialization, and request node checks. Since these RIs, presented in Table˜6, are a set of prompts, they can be modified and extended dynamically over time. In addition, CHF allows for manual modifications on the dialogue flow (Section˜A.3) and manual DST prompt optimization (Section˜A.4). We also attempt to refine the generated DST prompts ($P^{(s)}$) through automatic prompt optimization. Please refer to Section˜A.4 for details on automatic DST prompt optimization.
@@ 4 Experimental Settings
We use GPT-4o, GPT-5 (with reasoning levels of minimal and low), Claude 3.5 Sonnet, Gemini 2.0 Flash, and DeepSeek V3 (DSV3) as $\text{LLM}_{\text{GCG}}$ and $\text{LLM}_{\text{A}}$. Larger models are used for code generation—given the complexity of the task, we found that smaller models often fail to fully adhere to instructions. For further details, please refer to Section˜A.5.
Model Int.
Graph
Transfer
STAR MultiWOZ 2.2
F1 Accuracy BLEU JGA Inform Success BLEU Combined
SOLOIST - - - 35.9 81.7 67.1 13.6 88.0
MARS - - - 35.5 88.9 78.0 19.6 103.0
DARD - - - - 96.6 88.3 12.1 104.6
IG-TOD (few-shot) - - - 27 - 44 6.8 -
(Strict) Zero-shot Transfer
IG-TOD (zero-shot) - - - 13 - 31 4.2 -
BERT + Schema 29.7* 32.4* - - - - - -
SAM 51.2* 49.8* - - - - - -
AnyTOD xxl 68.0* 68.0* 44.3* 30.8 76.9 47.6 3.4 65.6
SGP-TOD 53.5 53.2 - - 82.0 72.5 9.2 86.5
$\text{CoDial}_{\text{free}}$
CoDial (4o, 4o-mini) - ri 36.6 36.1 23.0 - - - - -
$\text{CoDial}_{\text{structured}}$
CoDial (4o, 4o-mini) 58.5 60.1 45.2 28.4 76.6 54.6 3.5 69.1
CoDial (4o, 5-mini:l) 59.2 60.2 46.5 37.0 79.6 70.8 4.3 79.5
Table 1: Comparison of CoDial with baselines on STAR and MultiWOZ benchmarks. In “Strict Zero-Shot” the models have not seen a same task schema architecture in the training data. Results with an asterisk (*) are evaluated in a more relaxed, non-strict setting, and therefore, are not directly comparable. “Int.” stands for “Interpretable.” SAM results are cited from Zhao et al. (2023).
@@ 4.1 Datasets
The STAR dataset (Mosig et al., 2020), collected in a Wizard-of-Oz setup (human-human conversations), provides explicit task schemas (i.e., dialogue flows) to ensure consistent and deterministic system actions. We also use silver state annotations created in STARv2 (Zhao et al., 2023) for ablation studies. Refer to Section˜A.3 for more implementation details on STAR.
@@
MultiWOZ (Budzianowski et al., 2018) is a large-scale, multi-domain TOD dataset consisting of human-human conversations, with most domains involving booking subtasks such as hotel reservations and taxi services. Given the impracticality of crafting dialogue flows for every possible domain combination (Zhang et al., 2023), we report results in a naive oracle domain setting. Please refer to Section˜A.4 for more details.
@@ 4.2 Metrics
For the STAR dataset, we compute BLEU-4 score (Papineni et al., 2002) and follow Mosig et al. (2020) to compute F1 and accuracy. For the MultiWOZ dataset, we compute BLEU, Inform and Success rates, and Joint Goal Accuracy (JGA) using the official evaluation script (Nekvinda and Dušek, 2021). We report the mean of three runs. Refer to Section˜A.5 for more implementation details.
@@ 4.3 Baselines
For a complete list of compared methods, please refer to Section˜A.6. Our most comparable baselines are as follows:
@@
  • SGP-TOD (Zhang et al., 2023) is a zero-shot prompting approach that employs graph-based dialogue flows. Refer to Section˜A.6 for details on fair comparison.
For the remainder of this paper, by CoDial we refer to $\text{CoDial}_{\text{structured}}$ with GPT-4o and GPT-4o-mini as $\text{LLM}_{\text{GCG}}$ and $\text{LLM}_{\text{A}}$, respectively, unless otherwise specified.
@@ 5 Experimental Results
Superior Performance with Explicit Schemas
Table˜1 summarizes CoDial results on the benchmark datasets. CoDial achieves strong performance, surpassing all baselines except AnyTOD, and sets the new SOTA in strict zero-shot setting, where no same-architecture task schema is seen by the model. Our framework improves F1 by +5.7 and accuracy by +7 points over the previous SOTA. While AnyTOD achieves higher scores, it is evaluated in the easier non-strict setting and requires the user to write code, limiting accessibility to non-programmers. In contrast, CoDial operates in a graph-based transfer manner, eliminating the need for manual programming. We also observe that $\text{CoDial}_{\text{free}}$ lags behind $\text{CoDial}_{\text{structured}}$ and most baselines, indicating that LLMs struggle with unsupervised guardrailing code generation, likely due to limited availability of guardrailing languages, and still require human supervision.
Competitive Performance on MultiWOZ
Table˜1 also shows our results on the MultiWOZ dataset. Unlike STAR, where wizards were provided structured guidance for system responses, MultiWOZ lacks a predefined dialogue flow, making interactions less consistent. This variability in MultiWOZ poses additional challenges for heuristics-grounded and programmatic approaches like CoDial and AnyTOD. Consequently, CoDial is less effective on MultiWOZ. To address this, we experiment with GPT-5 with built-in reasoning as $\text{LLM}_{\text{A}}$ to improve the DST performance. We observe that CoDial achieves competitive performance with SOTA on Inform and Success metrics under the strict zero-shot setting, while maintaining interpretability. We further analyze the effect of DST performance in Section˜5.2. Similar to AnyTOD, CoDial relies on template-based outputs, which accounts for its lower BLEU score.
Model $\text{LLM}_{\text{GCG}}$ $\text{LLM}_{\text{A}}$ RI F1 Acc. BLEU
$\text{CoDial}_{\text{free}}$
CoDial - ri 4o 4o-mini 36.6 36.1 23.0
CoDial - ri Sonnet 4o-mini 32.1 31.8 18.0
$\text{CoDial}_{\text{structured}}$
CoDial original DFs 4o 4o-mini 51.9 51.1 38.9
CoDial - ri 4o 4o-mini 56.1 57.3 38.4
CoDial - ri Sonnet 4o-mini 57.0 58.4 39.2
CoDial DSV3 4o-mini 46.1 48.0 28.0
CoDial* Gem. 2 Fl. 4o-mini 50.5 52.1 32.9
CoDial Sonnet 4o-mini 57.7 58.5 39.3
CoDial 4o DSV3 55.6 56.8 44.2
CoDial 4o 4o-mini 58.5 60.1 45.2
CoDial 4o 5-mini:m 59.0 60.4 45.2
CoDial 4o 5-mini:l 59.2 60.2 46.5
@@
Table 2: Comparison of CoDial performance across different settings and ($\text{LLM}_{\text{GCG}}$, $\text{LLM}_{\text{A}}$) pairs on STAR dataset. The generated code for the model with an asterisk (*) has been manually fixed and is not directly comparable. DF stands for “dialogue flow.”
5.1 Detailed Analysis
Impact of Model Selection and CHF
@@ 5.1 Detailed Analysis
We experiment with different model choices for the ($\text{LLM}_{\text{GCG}}$, $\text{LLM}_{\text{A}}$) pairing (Table˜2). Better instruction following and more robust code generation often translate to higher overall performance. Because most LLMs are unfamiliar with guardrailing languages such as Colang, they must accurately interpret the $\text{prompt}_{\text{GCG}}$ to produce syntactically correct code. When the chosen LLM struggles with instruction following, code generation can fail, leading to incorrect or incomplete programs. Among the tested configurations, CoDial (4o, 5-mini) with built-in reasoning achieve the highest performance in all metrics. CoDial (4o, 4o-mini) performs comparably with lower cost and latency. Therefore, we use GPT-4o-mini for our ablations in Section˜5.2. We also report results in an oracle voting setting (Table˜5) between GPT-4o-mini and DSV3 as $\text{LLM}_{\text{A}}$, where for each task, we take the best-performing $\text{LLM}_{\text{A}}$ by F1. This results in an increase of +1.7+1.7 F1 and +1.5+1.5 accuracy.
Additionally, without modifications (Section˜A.3), the original STAR dialogue flows result in lower performance (F1: $51.9$). Manually modifying the dialogue flow and applying RIs to the generated code significantly enhances performance. We further explore the impact of LLM-aided corrections in Section˜5.2.
@@
Table 5: Ablations on MultiWOZ. Settings with an asterisk (*) are not directly comparable due to a simpler task setup.
Action Prediction and API Calls
We find that NeMo Guardrails’ intent detection performs strongly, achieving an F1 score of 96.3 on global actions (Table˜5). Additionally, we observe that STAR’s API calling precision—measured as the ratio of correct API calls to the total number of API calls—stands at 74.9. Table˜5 also summarizes the performance of the actions that are generated by $\text{LLM}_{\text{A}}$ (i.e., when NAP component does not generate an output). LLM-generated actions account for $25\%$ of all predicted actions, with $70\%$ of them belonging to three fallback actions: goodbye, out_of_scope, and anything_else. Excluding fallbacks, LLM-generated actions only account for $9.2\%$ of predictions, indicating that our NAP logic is generally effective at generating outputs based on the predicted state. Since fallback actions are a simple 3-way classification, we would expect high performance. However, $\text{LLM}_{\text{A}}$ achieves an F1 score of only $51.4$. We attribute this to the lack of an explicit schema for fallback actions in the STAR dataset, leading to inconsistencies in wizard annotations. Additionally, we observe a significant performance drop from fallback to non-fallback actions in both F1 ($51.4\rightarrow 38.7$) and accuracy. This suggests that despite having an explicit schema, LLMs struggle to capture the more complex logic needed to predict non-fallback actions. Our findings align with Dong et al. (2024), reinforcing the need for a neuro-symbolic approach.
State Prediction
Figure˜2 shows the error rate of the predicted conversation state across different node types for each model. To approximate the error, we compare the model’s predicted state with the estimated ground-truth state (i.e., the wizard’s state), as described in Section˜A.3. We find that the error rate generally inversely correlates with the overall performance in Table˜2; higher-performing models tend to exhibit lower state prediction error.
Single- vs. Multi-Domain Performance
Most of the MultiWOZ test set consists of multi-domain conversations, where a user may, for example, book both a taxi and a restaurant in the same dialogue. Since CoDial is designed for single-domain interactions, we report its performance on single-domain dialogues in Table˜5, where it performs well. However, with our naive oracle domain setting, CoDial performance drops significantly. This is likely due to compounded errors from DST to NAP, which we analyze further in Section˜5.2.
5.2 Ablation Studies
@@
To assess the impact of DST error, we evaluate CoDial under an Oracle setting. Since STAR does not provide gold DST labels, we simulate an oracle setting by using the silver annotations from STARv2 (Table˜5). This results in a performance gain of +2.2+2.2 F1 and +2.8+2.8 accuracy. We do the same for MultiWOZ, where we use the gold belief states, which leads to a substantial performance improvement (Table˜5). These findings suggest that investigating more advanced DST approaches, such as inference-time scaling explored in Section˜5, could be a promising direction to improve performance. We also experiment briefly with prompt optimization for lower cost DST improvements, described below.
@@
We use LLMs to perform iterative code refinement and automatic prompt optimization for the DST prompts. Refining the code with RIs consistently enhances CoDial’s performance, demonstrating the benefits of integrating user feedback into the generation process. After prompting the LLM to iteratively refine its outputs, CoDial achieves better accuracy and fluency (compared to CoDial - ri in Table˜2). We also conduct an ablation study to examine the effect of the individual RIs, summarized in Table˜5. Although all RIs are beneficial, most of the performance improvements can be attributed to the third RI, which refines the conditional logic of request nodes.
After observing the results of the oracle DST setting, we also apply prompt optimization to improve DST accuracy. As shown in Table˜5, automatic prompt optimization yields only marginal gains across metrics, with the exception of Inform, suggesting that automatic DST improvement remains a non-trivial challenge. To explore the impact of human feedback, we also experiment with manual prompt optimization (Section˜A.4), making minor edits to the prompts for the “attraction” domain. This results in consistent improvements across all metrics, reinforcing that human-crafted prompts can still outperform automatic optimization.
@@
To better understand the effectiveness of the proposed CoDial architecture, we experiment with a setting in which the NAP component is removed and all actions are predicted in a fully generative manner by the $\text{LLM}_{\text{A}}$, similar to Zhang et al. (2023). We prompt the $\text{LLM}_{\text{A}}$ with simplified dialogue flows following their work and include predicted DST slots in the prompt. This results in a substantial drop in performance (Table˜5), highlighting the importance of our NAP approach. We further ablate the model by removing the DST component, causing a larger drop as expected.